OpenStudy (anonymous):

help!! F(X)=85+15X B.What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain.

OpenStudy (kirbykirby):

The domain is the set of possible values for which this function would "work" (that is there are no tricky "divided by 0" or similar aspects to worry about). Now, what values work for F(X)? That is, what values of "X" can you put in F(X) and it will still give you a value?

OpenStudy (anonymous):

all whole numbers

OpenStudy (kirbykirby):

Would the value 1.6 work?

OpenStudy (anonymous):

all numbers

OpenStudy (kirbykirby):

Unfortunately in math we need to be precise. What do you mean by "all numbers". (Natural numbers, Integers, etc.?)

OpenStudy (anonymous):

all real numbers

OpenStudy (kirbykirby):

yes indeed.

OpenStudy (anonymous):

there is a scenario that goes along with it can I send it to you?

OpenStudy (kirbykirby):

ok

OpenStudy (anonymous):

1.Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15.

OpenStudy (kirbykirby):

Oh there's a scenario... ok well then the domain might not be all real numbers...

OpenStudy (kirbykirby):

Yes indeed this should have been posted before asking F(X)=85+15X's domain. Abstractly, this function can take any real number. But now we are considering a real physical situation. So now X would be the number of charms for a bracelet. So the domain would be \(X \in\{0,1,2,3,...\}\). Theoretically it can go to infinity (even though it's physically impossible on a bracelet), but we have no further information.

OpenStudy (kirbykirby):

So the domain is the set of non-negative integers, which can also be written as \(\mathbb{Z}^+ \cup \{0\}\)

OpenStudy (anonymous):

why cant it be 1.5 ?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

can you put 1 and a half charms on a bracelet?

OpenStudy (anonymous):

no that was a dumb question sorry

OpenStudy (kirbykirby):

No hehe don't worry :P no dumb question

OpenStudy (anonymous):

@kirbykirby can I put positive whole numbers

OpenStudy (kirbykirby):

The term "whole number" is sometimes not clearly defined (if it includes "0" or not). If it's properly defined in your course, then you can use it. Now I would say though that you should include "0" because you could very well just buy the bracelet on its own, without any charms. (Recall X = the number of charms, if X =0, then you get F(0) = 85 +15(0) = 85, the cost of the bracelet by itself).

OpenStudy (anonymous):

so positive whole numbers will be fine

OpenStudy (kirbykirby):

no, positive numbers excludes 0. You should use the term "non-negative" instead of "positive". Non-negative means all positive numbers AND zero.

OpenStudy (anonymous):

C.If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250.

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):



OpenStudy (kirbykirby):

each price on the y-axis is increasing by 15, since adding a new charm adds $15. The x-axis is labelling the # of charms. So at X =4, F(X) = 165 for example.

OpenStudy (kirbykirby):

There should be another point at X =9, F(X) = 240 (can't exceed that amount since the next amount would be 255, but you only have $250)

OpenStudy (kirbykirby):

urhg sorry I missed up on the first step... I added 25 instead of 15 oye. So the axis should be going as: 85, 100, 115, 130, 145, 160, 175, 190, 205, 220, 235, 250

OpenStudy (anonymous):

okay yes I noticed that

OpenStudy (anonymous):

do I connect them?

OpenStudy (kirbykirby):

No you shouldn't. You have the exact values for your function. Remember, we specified the domain as the non-negative whole numbers.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

E.As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench. Price for bracelet 100$ each charm is 12$. Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers.

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

can you help me on the question I just asked